# The dimension and Bose distance of some BCH codes of length $\frac{q^{m}-1}{\lambda}$

### v1

@@ VIII Some non-narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta,b)}$
Let $m\geq 4$ and $b$ be two integers with $\frac{q^{m-h}-1}{\lambda}+1\leq b<\frac{q^{\lfloor({2m-1})/{3}\rfloor+1}-1}{\lambda}$. Let $j_{\delta}=\lfloor\log_{q}(\lambda\delta)\rfloor-h$, let $\sum\limits_{\ell=0}^{h+j_{b}}b_{\ell}q^{\ell}$ be the $q$-adic expansion of $\lambda b$, and let $r_{b}$ be the smallest integer in $m-2h-j_{b},j_{b}]$ such that $b_{h+r_{b}}>0$. If $\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{b}}^{h+j_{b}}b_{\ell}q^{\ell-(h-r_{b})}$, then
$\mathrm{dim}(\mathcal{C}_{(q,n,\delta,b)})=n-m(\delta-1)$ (60)
@@
$d_{B}(\mathcal{C}_{(q,n,\delta,b)})=\delta$ (61)
for any integer $\delta$ such that $2\leq\delta\leq\frac{q^{h-j_{b}}-1}{\lambda}-\lfloor\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rfloor+1$.
@@
where $t_{a}$ is largest integer such that $q^{t_{a}}\mid a$. We now show that
@@
Case 1. Suppose that $t_{a}=0$, i.e., $q\nmid a$. Then $f(a)=a$. It follows that $V(\lambda f(a))$ has the form given in (62). By Remarks 4 and 5, this implies that $\lambda f(a)\not\in\mathcal{B}_{j_{b}}(i)$ for any $i\in[-j_{b},j_{b}]$. Applying Theorem 1, we conclude that $\lambda f(a)\not\in\mathcal{H}\cup\mathcal{S}$. Therefore, $\lambda f(a)\in\mathcal{L}_{m}^{\lambda n}$. By Lemma 1, this implies that $f(a)\in\mathcal{L}^{n}_{m}$.
@@
Furthermore, the inequality $\sum\limits_{\ell=0}^{h-j_{b}-1}a_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{b}}^{h+j_{b}}b_{\ell}q^{\ell-(h-r_{b})}$ implies that $a_{j_{b}+r_{b}}>0$. It follows that $\lambda f(a)\not\in\mathcal{B}_{j_{b}-t_{a}}(i)$ for any $i\in[t_{a}-j_{b},j_{b}-t_{a}]$, and hence $\lambda f(a)\not\in\mathcal{H}\cup\mathcal{S}$. This implies that $\lambda f(a)\in\mathcal{L}^{\lambda n}_{m}$, and hence $f(a)\in\mathcal{L}^{n}_{m}$.
Case 3. Suppose $t_{a}\geq j_{b}+1$. In this case, we have $\lambda f(a)<q^{h}$ and $\lambda f(a)q^{h}\bmod\lambda n\neq\lambda f(a)$. By Lemma 2, we have $|C_{\lambda n}(\lambda f(a))|=m$. Moreover, by applying [35, Theorem 2.3], we conclude that $\lambda f(a)$ is a coset leader modulo $\lambda n$. It follows that $f(a)\in\mathcal{L}^{n}_{m}$.
By now we have demonstrated that (63) holds. Noticing that $C_{n}(a)=C_{n}(f(a))$ and $f$ is injective, it follows that
@@
for any integer $\delta$ such that $2\leq\delta\leq\frac{q^{h-j_{b}}-1}{\lambda}-\lfloor\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rfloor+1$. Therefore, both (60) and (61) hold. ∎
@@
Applying the above theorem, we find the following optimal BCH codes.
@@ IX Conclusion
In this paper, we investigate the dimension and Bose distance of BCH codes of length $\frac{q^{m}-1}{\lambda}$, where $\lambda$ is a positive divisor of q1q-1. Our main contribution is to provide explicit formulas for the dimension and the narrow-sense Bose distance of BCH codes of length $\frac{q^{m}-1}{\lambda}$ for a much larger range of designed distances than previously known. In addition, we extend these results to some non-narrow-sense BCH codes of length $\frac{q^{m}-1}{\lambda}$. Applying our results, we find some BCH codes with good parameters.
@@ Appendix H Proof of Lemma 10
The arguments used to derive (11) and (12) are analogus, so we only demonstrate equation (11) holds through the following two cases.
@@
$\sum\limits_{t\in\mathcal{W}_{i}}\frac{tq^{2i-2}(q-1)}{\lambda}$ $=$ $\sum\limits_{j=0}^{q^{k-i-1}(a-1)-1}\frac{(q^{k-i}+qj)q^{2i-2}(q-1)}{\lambda}$ (75)
$=$ $\frac{1}{2\lambda}(q-1)(a-1)\left[(a+1)q^{2k-3}+q^{k+i-2}\right]$
@@ Appendix I Proof of Assertion 1
Recalling the definition of $s_{\delta}$, it follows that $s_{\delta}$ is the smallest integer in $k_{\delta}+1,k_{\delta}]$ such that $a_{h+s_{\delta}}>0$. By Remark 5, this implies $a\not\in\mathcal{A}_{k_{\delta}}(i)$ for any integer $i\neq s_{\delta}$. Therefore, $\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(i)=\varnothing$ for any integer $i\neq s_{\delta}.$ Then applying Theorem 1, we can conclude that (79) holds.
Now, let us count the number of integers in the set $\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(s_{\delta})\cap\mathcal{D}_{\lambda}.$ Recall that $w_{\delta}=\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$, $\mu(\delta)=\min\left\{\sum\limits_{\ell=0}^{h-k_{\delta}}\delta_{\ell}q^{\ell},\sum\limits_{\ell=h-s_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell-(h-s_{\delta}+1)}\right\}$, and $\alpha(a)=\sum\limits_{\ell=0}^{k_{\delta}+s_{\delta}-1}a_{\ell}q^{\ell}$ for each integer $a\in\mathcal{A}_{k_{\delta}}(s_{\delta})$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k_{\delta}}a_{\ell}q^{\ell}$. We conclude from equation (80) and the definition of $\mathcal{A}_{k_{\delta}}(s_{\delta})$ that an integer $a\in\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(s_{\delta})\cap\mathcal{D}_{\lambda}$ if and only if $a$ admits the decomposition
@@
Suppose that $m$ is even. The equality in (20) can be obtained by emplying a similar argument as above. It remains to establish equation (21). By applying Lemma 2, we can conclude that $a\in\left[\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}$ if and only if $V(a)$ has the form
@@
Then applying Theorem 1, it follows that (22) holds.
@@
Then applying Lemmas 7 and 8, the assertion follows. ∎
@@
By applying Lemma 2, we can conclude that an integer $a\in\mathcal{H}\cap[q^{h+k},q^{h+k+1})\cap\mathcal{D}_{\lambda}$ if and only if $a$ adimits the decomposition

### v2

@@ VIII Some non-narrow-sense BCH codes of length $(q^{m}-1)/\lambda$
Let $m\geq 4$ be an integer, and let $\lambda$ be a positive divisor of q1q-1. Set $n=(q^{m}-1)/\lambda$ and $h=\lfloor\frac{m}{2}\rfloor$. Let $b$ be an integer with $\frac{q^{m-h}-1}{\lambda}+1\leq b<\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}.$ Define $j_{b}=\lfloor\log_{q}(\lambda b)\rfloor-h,$ and write the $q$-adic expansion of $\lambda b$ as $\lambda b=\sum\limits_{\ell=0}^{h+j_{b}}b_{\ell}q^{\ell}.$ Let $r_{b}$ be the smallest integer in $m-2h-j_{b},\;j_{b}\,]$ such that $b_{h+r_{b}}>0$. If $\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}\;>\;\sum\limits_{\ell=h-r_{b}}^{h+j_{b}}b_{\ell}q^{\ell-(h-r_{b})},$ then for any integer $\delta$ with $2\leq\delta\leq\frac{q^{h-j_{b}}-1}{\lambda}-\left\lfloor\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}/\lambda\right\rfloor+1,$ we have
$\dim(\mathcal{C}_{(q,n,\delta,b)})=n-m(\delta-1)$ (62)
@@
$d_{B}(\mathcal{C}_{(q,n,\delta,b)})=\delta.$ (63)
@@
where $t_{a}$ is largest integer such that $q^{t_{a}}\mid a$. Recall that $\mathcal{L}^{n}_{\ell}(0,n-1)$ is simply denoted by $\mathcal{L}^{n}_{\ell}$. We now show that
@@
Case 1. Suppose that $t_{a}=0$, i.e., $q\nmid a$. Then $f(a)=a$. It follows that $V(\lambda f(a))$ has the form given in (64). By Remarks 4 and 5, this implies that $\lambda f(a)\not\in\mathcal{B}_{j_{b}}(i)$ for any $i\in[-j_{b},j_{b}]$. Applying Lemma 3, we conclude that $\lambda f(a)\not\in\mathcal{H}\cup\mathcal{S}$. Therefore, $\lambda f(a)\in\mathcal{L}_{m}^{\lambda n}$. By Lemma 1, this implies that $f(a)\in\mathcal{L}^{n}_{m}$.
@@
Furthermore, the inequality $\sum\limits_{\ell=0}^{h-j_{b}-1}a_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{b}}^{h+j_{b}}b_{\ell}q^{\ell-(h-r_{b})}$ implies that $a_{j_{b}+r_{b}}>0$. It follows that $\lambda f(a)\not\in\mathcal{B}_{j_{b}-t_{a}}(i)$ for every integer $i\in[t_{a}-j_{b},j_{b}-t_{a}]$, and hence $\lambda f(a)\not\in\mathcal{H}\cup\mathcal{S}$. This implies that $\lambda f(a)\in\mathcal{L}^{\lambda n}_{m}$, and hence $f(a)\in\mathcal{L}^{n}_{m}$.
Case 3. Suppose $t_{a}\geq j_{b}+1$. In this case, we have $\lambda f(a)<q^{h}$ and $\lambda f(a)q^{h}\bmod\lambda n\neq\lambda f(a)$. By Lemma 4, we have $|C_{\lambda n}(\lambda f(a))|=m$. Moreover, by applying Lemma 2, we conclude that $\lambda f(a)$ is a coset leader modulo $\lambda n$. It follows that $f(a)\in\mathcal{L}^{n}_{m}$.
By now we have demonstrated that (65) holds. Note that $C_{n}(a)=C_{n}(f(a))$ and $f$ is injective. Therefore, we have
@@
for any integer $\delta$ such that $2\leq\delta\leq\frac{q^{h-j_{b}}-1}{\lambda}-\lfloor\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rfloor+1$. Recalling equation (2) and the fact that the Bose distance $d_{B}$ of $\mathcal{C}_{(q,n,\delta,b)}$ is the largest integer for which (3) holds, we deduce that (62) and (63) both hold. ∎
Applying the above theorem, we find the following BCH codes.
@@
By the Hamming bound, no linear code over $\mathbb{F}_{3}$ with parameters $[80,76,d]$ exists for $d>32$, nor does a linear code over $\mathbb{F}_{4}$ with parameters $[255,251,d]$ exist for $d>2$. Therefore, the BCH codes in the above examples are optimal. Readers can also find the bounds for the minimum distance of linear codes of length up to 256 in the Database.
@@ IX Conclusion
In this paper, we investigate the dimension and Bose distance of BCH codes of length $(q^{m}-1)/\lambda$, where $\lambda$ is a positive divisor of q1q-1. Our main contribution is to provide explicit formulas for the dimension and Bose distance of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$ for a much larger range of designed distances than previously known. In addition, we extend these results to some non-narrow-sense BCH codes of the same length. Applying our results, we find some BCH codes with good parameters.
@@ Appendix H Proof of Lemma 13
The arguments used to derive (13) and (14) are analogous, so we only demonstrate equation (13) holds through the following two cases.
@@
$\begin{split}\sum\limits_{t\in\mathcal{W}_{i}}\frac{tq^{2i-2}(q-1)}{\lambda}&=\sum\limits_{j=0}^{q^{k-i-1}(a-1)-1}\frac{(q^{k-i}+qj)q^{2i-2}(q-1)}{\lambda}\\ &=\frac{1}{2\lambda}(q-1)(a-1)\left[(a+1)q^{2k-3}+q^{k+i-2}\right]\end{split}$ (77)
@@ Appendix I Proof of Assertion 1
Recalling the definition of $s_{\delta}$, it follows that $s_{\delta}$ is the smallest integer in $k_{\delta}+1,k_{\delta}]$ such that $a_{h+s_{\delta}}>0$. By Remark 5, this implies $a\not\in\mathcal{A}_{k_{\delta}}(i)$ for any integer $i\neq s_{\delta}$. Therefore, $\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(i)=\varnothing$ for any integer $i\neq s_{\delta}.$ Then applying Lemma 3, we can conclude that (81) holds.
Now, let us count the number of integers in the set $\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(s_{\delta})\cap\mathcal{D}_{\lambda}.$ Recall that $w_{\delta}=\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$, $\mu(\delta)=\min\left\{\sum\limits_{\ell=0}^{h-k_{\delta}}\delta_{\ell}q^{\ell},\sum\limits_{\ell=h-s_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell-(h-s_{\delta}+1)}\right\}$, and $\alpha(a)=\sum\limits_{\ell=0}^{k_{\delta}+s_{\delta}-1}a_{\ell}q^{\ell}$ for each integer $a\in\mathcal{A}_{k_{\delta}}(s_{\delta})$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k_{\delta}}a_{\ell}q^{\ell}$. We conclude from equation (82) and the definition of $\mathcal{A}_{k_{\delta}}(s_{\delta})$ that an integer $a\in\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(s_{\delta})\cap\mathcal{D}_{\lambda}$ if and only if $a$ satisfies the decomposition
@@
Suppose that $m$ is even. The equality in (22) can be obtained by employing a similar argument as above. It remains to establish equation (23). By applying Lemma 5, we can conclude that $a\in\left[\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}$ if and only if $V(a)$ has the form
@@
We may now conclude that, for each integer $i\in[-k_{\delta}+1,\,k_{\delta}]$, an integer
$a\in\left[q^{h+k_{\delta}},\sum_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{A}_{k_{\delta}}(i)\cap\mathcal{D}_{\lambda}$
exists if and only if $a$ admits a decomposition of the form
$a=t(a)\,q^{h+i}+\alpha(a),$
where
$\begin{cases}\lambda\mid t(a)+\alpha(a),\\ t(a)\in\mathcal{T}_{i}(\delta),\\ 1\leq\alpha(a)\leq\lfloor t(a)\,q^{2i-1}\rfloor\quad\text{and}\quad q\nmid\alpha(a).\end{cases}$
@@
Then applying Lemma 3, it follows that (24) holds.
@@
Then applying Lemmas 10 and $1$, the assertion follows. ∎
@@
By applying Lemma 5, we can conclude that an integer $a\in\mathcal{H}\cap[q^{h+k},q^{h+k+1})\cap\mathcal{D}_{\lambda}$ if and only if $a$ admits the decomposition